OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 4 degrees Celsius is.

OpenStudy (anonymous):

part 1 F= (c/(5/9))+32 part 2 Fahrenheit 4 = -15.55 C

OpenStudy (anonymous):

thxx :)

OpenStudy (anonymous):

how can i solve the first part?

OpenStudy (anonymous):

C=5/9(F-32) Fe need to solve for F. Since F is in Parentheses first divide (5/9) on both sides. We get C/(5/9)=F-32 Now Add (-32) to both sides. since it's being subtracted to F we get (C/(5/9))+32)=F F=(C/(5/9))+32) OR Which is same thing F=32+(C/(5/9))

OpenStudy (anonymous):

thxx